(ii) Let a be a number of the form a = 6n + 1 or a = 6n−1 for some integer n. Show that a2 ≡ 1 (mod 24). (Hint: Calculate (6n±1)2 and consider two different cases for n.)
Answer #1

If a=6n+1 then a^2=36n^2+12n+1 . If a=6n-1 then a^2=36n^2-12n+1 . Now, there are two cases:

Case I: n is even, say n=2m ; then either

a^2=36n^2+12n+1=144m^2+24m+1=24(6m^2+m)+1

or

a^2=36n^2-12n+1=144m^2-24m+1=24(6m^2-m)+1

Thus, either way we have a^2\equiv 1\mod 24 .

Case II: n is odd, say n=2m+1 ; then either

\begin{align*}a^2&=36n^2+12n+1\\ &=144m^2+144m+36+24m+12+1\\ &=24(6m^2+7m+2)+1\end{align*}

or

\begin{align*}a^2&=36n^2-12n+1\\ &=144m^2+144m+36-24m-12+1\\ &=24(6m^2+5m+1)+1\end{align*}

Thus, either way we have a^2\equiv 1\mod 24 .

This proves that a^2\equiv 1\mod 24 ​​​​​​​.